# Recent Advances of Foundation Language Models-based Continual Learning: A Survey

### v1

@@ 4.1.3. VLMs-based DIL
Vision-language models (VLMs) have demonstrated superiority in domain-incremental learning contexts. Yi et al. (Yi et al., 2023) integrate VLMS with continual learning methodologies to develop a general-purpose medical AI. Moreover, their study highlights the significance of data-efficient adaptation algorithms in minimizing the necessity for extensive labeling when transitioning to new domains or tasks. Furthermore, the prompt text is utilized to master the pre-trained knowledge embedded within VLMs. Aiming to independently learn prompts across disparate domains by using pre-trained VLMs, S-Prompt (Wang et al., 2022b) is devised (Figure 5(c)). This method encompasses techniques for acquiring image prompts and introduces an innovative methodology for language-image prompt acquisition. Prompt learning is conducted separately, utilizing a unified cross-entropy loss function during training. During inference, a K-NN (k-nearest neighbors) technique is employed to discern the domain.
In the domain of visual question answering, Zhang et al. (Zhang et al., 2023d) introduce VQACL, a novel framework designed to effectively integrate data from both visual and linguistic modalities. This integration is achieved through a dual-level task sequence that enhances the model’s performance on complex multimodal tasks. Central to VQACL is a compositionality test that evaluates the model’s ability to generalize new skill and concept combinations. The framework also incorporates a novel representation learning strategy that differentiates between sample-specific (SS) and sample-invariant (SI) features. SS features capture distinctive attributes of individual inputs, enhancing output uniqueness, while SI features, derived from category prototypes, ensure essential characteristics are retained.
@@
4.2. Task-Incremental Learning
4.2.1. PLMs-based TIL
@@
Drawing inspiration from neurobiological mechanisms, Maekawa et al. (Maekawa et al., 2023) present an inventive approach known as Hippocampal Memory Indexing (HMI) to augment the generative replay technique. HMI leverages hippocampal memory indexing to integrate compressed representations of prior training instances, facilitating selective guidance for the generation of training samples. This methodological refinement contributes to heightened specificity, balance, and overall quality of the replayed samples. In tackling the Continual Few-Shot Relation Learning (CFRL) challenge, Qin et al. (Qin and Joty, 2022a) propose ERDA as a solution, drawing upon replay- and regularization-based techniques. The ERDA framework integrates embedding space regularization and data augmentation strategies to effectively confront the task of acquiring new relational patterns from scarce labeled instances while mitigating the risk of catastrophic forgetting. Wang et al. (Wang et al., 2019b) introduced a memory-based approach to continual learning, termed Episodic Memory Replay (EMR). This method leverages working memory by selectively replaying stored samples during each iteration of learning new tasks, thereby facilitating the integration of new knowledge while preserving previously acquired information.
Conure (Yuan et al., 2021) is a framework that effectively manages multiple tasks by leveraging the redundancy of parameters in deep user representation models. Initially, it prunes less critical parameters to make room for new, task-specific ones. Subsequently, it incorporates these new parameters while retaining key parameters from prior tasks, facilitating positive transfer learning. To prevent the loss of previously acquired knowledge, it maintains these essential parameters in a fixed state. Notably, TERACON (Kim et al., 2023) utilizes task-specific soft masks to isolate parameters, which not only targets parameter updates during training but also clarifies the relationships between tasks. This method includes a novel knowledge retention module that utilizes pseudo-labeling to mitigate the well-known problem of catastrophic forgetting.
Ke et al. (Ke et al., 2021a) introduced a model known as CTR, which employs innovative techniques such as CL-plugin and task masking to tackle the issue of catastrophic forgetting and to facilitate knowledge transfer across tasks. These strategies are particularly effective when utilized in conjunction with pre-trained language models, such as BERT, enhancing their adaptability and efficacy. In the specific context of User Intent Classification (UIC) within large-scale industrial applications, Wang et al. (Wang et al., 2021a) introduce a novel methodology, MeLL, which utilizes a BERT-based text encoder to generate robust, dynamically updated text representations for continual learning. MeLL combines global and local memory networks to preserve prototype representations across tasks, acting as a meta-learner that rapidly adapts to new challenges. It employs a Least Recently Used (LRU) policy for efficient global memory management and minimizes parameter growth.
Recent advancements in conversational AI have concentrated on mitigating the limitations of traditional chatbots, which are dependent on static knowledge bases and extensive manual data annotation. The Lifelong INteractive learning in Conversation (LINC) methodology represents a significant stride in this direction, embodying a dynamic learning framework that mimics human cognitive processes during interactions (Liu and Mazumder, 2021; Mazumder and Liu, 2024; Liu, 2020). Structured around an Interaction Module, a Task Learner, and a Knowledge Store, LINC enables chatbots to dynamically integrate and utilize knowledge within conversational contexts (Figure LABEL:fig:LINC). This approach not only facilitates real-time information extraction and learning from user interactions but also addresses challenges such as managing erroneous inputs, refining conversational skills, and maintaining user engagement, thereby enhancing the chatbots’ linguistic and interactive capabilities.
Continual Pre-training Methods.
Continual pre-training represents a paradigm wherein pre-trained language models (PLMs) are progressively enhanced through the assimilation of new knowledge from expanding datasets. Sun et al. (Sun et al., 2020b) introduced a framework called ERNIE 2.0, which incrementally constructs pre-training tasks, enabling the acquisition of complex lexical, syntactic, and semantic nuances embedded within the training corpora. This model deviates from traditional fixed-task training, instead employing continual multi-task learning to refine its capabilities continually. Further advancing the field of continual pre-training, RecyclableTuning (Qin et al., 2023) introduces a novel concept of recyclable tuning that features two distinct strategies: initialization-based and distillation-based methods. The former uses fine-tuned weights of existing PLMs as the basis for further enhancements, capitalizing on the established parametric relationships. Conversely, the distillation-based approach harnesses outdated weights to maintain knowledge continuity and efficiency in successor models.
@@
Expounding upon the crucial need for more efficacious knowledge integration, Zhang et al. (Zhang et al., 2022b) propose a pioneering strategy that entails the integration of Adaptive Compositional Modules alongside a replay mechanism. These modules are designed to dynamically adjust to new tasks and are supplemented by pseudo-experience replay, significantly enhancing knowledge transfer. This framework is distinguished by its adaptive integration of modules within transformer architectures, skillfully orchestrating the interactions between existing and new modules to address emerging tasks. Additionally, the implementation of pseudo-experience replay promotes efficient knowledge transfer across these modules. Concurrently, Jin et al. (Jin et al., 2021a) introduce the Continual Learning of Few-Shot Learners (CLIF) challenge, wherein a model accumulates knowledge continuously across a series of NLP tasks. Their investigation delves into the impact of continual learning algorithms on generalization capabilities and advances a novel approach for generating regularized adapters.
@@
Zhao et al. (Zhao et al., 2022) propose the Prompt Conditioned VAE for Lifelong Learning (PCLL) specifically designed for task-oriented dialogue (ToD) systems. PCLL employs a conditional variational autoencoder influenced by natural language prompts to generate high-quality pseudo samples, effectively capturing task-specific distributions. Additionally, a distillation process is integrated to refine past knowledge by reducing noise within pseudo samples. Razdaibiedina et al. (Razdaibiedina et al., 2023) introduce Progressive Prompts (PP), a novel approach to continual learning in language models. PP addresses catastrophic forgetting without resorting to data replay or an excessive proliferation of task-specific parameters. This method involves acquiring a fresh soft prompt for each task, gradually appending it to previously learned prompts while keeping the base model unchanged. The ConTinTin paradigm (Yin et al., 2022) develops a computational framework for sequentially mastering a series of new tasks guided by explicit textual instructions. It synthesizes projected outcomes for new tasks based on instructions, while facilitating knowledge transfer from prior tasks (forward-transfer) and maintaining proficiency in previous tasks (backward-transfer).
In the realm of lifelong in-context instruction learning aimed at enhancing the target PLM’s instance- and task-level generalization performance as it observes more tasks, DYNAINST is devised by Mok et al. (Mok et al., 2023). DYNAINST integrates the principles of parameter regularization and experience replay. The regularization technique employed by DYNAINST is tailored to foster broad local minima within the target PLM. In order to devise a memory- and computation-efficient experience replay mechanism, they introduce Dynamic Instruction Replay, comprising Dynamic Instance Selection (DIS) and Dynamic Task Selection (DTS). DIS and DTS dynamically determine the selection of instances and tasks to be stored and replayed, respectively.
4.2.2. LLMs-based TIL
Recent attention has focused on the convergence of large language models (LLMs) with continual learning methodologies, exemplified by significant contributions such as those by Wang et al. (Wang et al., 2024a) and Peng et al. (PENG et al., 2024). Benefiting from vast corpora and advanced hardware infrastructure, LLMs showcase remarkable capabilities in language comprehension and generation. However, challenges arise in scenarios involving sequential tasks, where LLMs often exhibit a decline in performance known as catastrophic forgetting.
Traditional Methods.
DynaMind, introduced by Du et al. (Du et al., 2023), emerges as a pioneering framework that intricately incorporates memory mechanisms and modular operators, enhancing the precision of LLM outputs. Comprising three essential components, DynaMind includes a memory module dedicated to storing and updating acquired knowledge, a modular operator for processing input data, and a continual learning module responsible for dynamically adjusting LLM parameters in response to new knowledge. Furthermore, Luo et al. (Luo et al., 2023) conduct a thorough investigation into catastrophic forgetting (CF) in Large Language Models (LLMs) during continual fine-tuning. Their experiments across various domains, including domain knowledge, reasoning, and reading comprehension, reveal the prevalence of CF in LLMs ranging from 1b to 7b scale, with severity increasing with model size. Comparative analysis between decoder-only BLOOMZ and encoder-decoder mT0 indicates that BLOOMZ exhibits less forgetting. Additionally, LLMs demonstrate the ability to mitigate language bias during continual fine-tuning. Contrasting ALPACA against LLAMA, the study highlights ALPACA’s superiority in preserving knowledge and capacity, suggesting that general instruction tuning aids in mitigating CF during subsequent fine-tuning phases. This research provides valuable insights into CF dynamics in LLMs, offering strategies for knowledge retention and bias mitigation.
Wang et al. (Wang et al., 2024a) present TRACE, an innovative benchmark meticulously designed to evaluate continual learning capabilities in LLMs. Comprising eight distinct datasets spanning challenging tasks, including domain-specific challenges, multilingual capabilities, code generation, and mathematical reasoning, TRACE serves as a comprehensive evaluation platform. The authors rigorously examine the effectiveness of conventional Continual Learning (CL) methods when applied to LLMs within the TRACE framework. Peng et al. (PENG et al., 2024) propose the Joint Adaptive ReParameterization (JARe) framework, enhanced with Dynamic Task-related Knowledge Retrieval (DTKR), to facilitate adaptive adjustment of language models tailored to specific downstream tasks. This innovative approach leverages task distribution within the vector space, aiming to streamline and optimize the continual learning process seamlessly.
Parameter-Efficient Tuning Methods.
Large language models (LLMs) encounter several substantial challenges that limit their practical applications. These include high computational requirements, significant memory demands, and a tendency toward catastrophic forgetting. Such limitations highlight the need for ongoing research into more efficient and robust approaches to training and deploying these models. Continual Parameter-Efficient Tuning (ConPET) (Song et al., 2023) is designed for the continuous adaptation of LLMs across diverse tasks, leveraging parameter-efficient tuning (PET) strategies to enhance both efficiency and performance. ConPET encompasses two primary modes: Static ConPET and Dynamic ConPET. Static ConPET adapts techniques previously utilized in smaller models for LLMs, thus minimizing tuning costs and reducing the risk of overfitting. Conversely, Dynamic ConPET enhances scalability by employing distinct PET modules for varying tasks, supplemented by a sophisticated selector mechanism.
Moreover, the ELM strategy (Jang et al., 2023) involves initially training a compact expert adapter on the LLM for each specific task, followed by deploying a retrieval method to select the most appropriate expert LLM for each new task. Furthermore, Wang et al. (Wang et al., 2023a) have proposed orthogonal low-rank adaptation (O-LoRA), a straightforward yet efficacious method for facilitating continual learning in language models. O-LoRA mitigates catastrophic forgetting during task acquisition by employing distinct low-rank vector subspaces maintained orthogonally to minimize task interference. This method highlights the potential of orthogonal subspace techniques in improving the adaptability of language models to new tasks without compromising previously acquired knowledge.
Instruction Tuning-based Methods.
Scialom et al. (Scialom et al., 2022a) introduced Continual-T0, an innovative framework aimed at exploring the capabilities of large language models (LLMs) through continual learning, incorporating rehearsal techniques. A central aspect of this approach is the employment of instruction tuning, a key strategy designed to enhance the adaptability and effectiveness of LLMs when encountering novel tasks. Leveraging self-supervised pre-training, Continual-T0 demonstrates exceptional proficiency in mastering new language generation tasks while maintaining high performance across a diverse range of 70 previously encountered datasets. Despite the demonstrated proficiency of LLMs in adhering to instructions, their ability to generalize across underrepresented languages remains suboptimal. In response, InstructAlign (Cahyawijaya et al., 2023) is proposed to address this challenge by aligning newly introduced languages with those previously learned, which possess abundant linguistic resources, thereby mitigating instances of catastrophic forgetting. The core novelty of this approach lies in its advancement of language adaptation methodologies for instruction-tuned LLMs, with particular emphasis on integrating underrepresented languages.
@@ 4.2.3. VLMs-based TIL
CTP (Zhu et al., 2023) employs topology preservation and momentum contrast to maintain consistent relationships within sample mini-batches across tasks, thereby preserving the distribution of prior embeddings. CTP also introduces the P9D dataset, comprising over one million image-text pairs across nine domains, aimed at visual language continuous pre-training (VLCP). Zheng et al. (Zheng et al., 2023) address the issue of zero-shot transfer degradation in visual language models by introducing the Zero-Shot Continual Learning (ZSCL) method. This novel approach utilizes a label-free dataset to facilitate distillation in the feature space, coupled with the application of weight regularization within the parameter space. Furthermore, they introduce a new benchmark, the Multi-Domain Task Incremental Learning (MTIL), designed to evaluate incremental learning strategies across various domains, thereby enhancing the assessment of such methods. Moreover, ZSCL has also been adapted for use in the CIL setting, further broadening its applicability
@@
By decoupling prompts and prompt interaction strategies, TRIPLET (Qian et al., 2023) effectively captures complex interactions between modalities. This includes specific designs for visual, textual, and fused prompts, as well as how to interact between different tasks through these prompts and retain crucial information, thereby reducing catastrophic forgetting. Decoupled prompts are designed to separate prompts in terms of multi-modality, layer-wise, and complementary, with each type of prompt containing learnable parameters intended to capture modality-specific knowledge from pre-trained visual-language models and training data. The prompt interaction strategies consist of three main components: the Query-and-Match Strategy, the Modality-Interaction Strategy, and the Task-Interaction Strategy. These components work together to enhance the model’s adaptability to different tasks and its memory for old tasks.
Moreover, COIN (Chen et al., 2024) introduces a new continuous instruction tuning benchmark designed to evaluate the performance of Multimodal Large Language Models (MLLMs) in the sequential instruction fine-tuning paradigm. CoIN includes 10 commonly used data sets covering 8 task categories, ensuring the diversity of instructions and tasks. In addition, the trained model is evaluated from two aspects: instruction following and general knowledge, which respectively evaluate the consistency with human intention and the preservation of knowledge for reasoning. CoIN converts commonly used visual-linguistic datasets into instruction fine-tuning data formats by using different templates to explore the behavior of MLLMs in continuous instruction fine-tuning. This method takes into account the diversity between different tasks and attempts to enhance the model’s adaptability to new and old tasks through diversified instruction templates. To alleviate the catastrophic forgetting problem of MLLMs, CoIN introduces the MoELoRA method. This method reduces forgetting by using different experts to learn knowledge on different tasks and using gate functions to regulate the output of these experts.
Parameter-Efficient Methods.
MoE-Adapters4CL (Yu et al., 2024) introduces a parameter-efficient continual learning method to mitigate long-term forgetting in incremental learning of visual-language models. Their approach involves dynamically extending the pre-trained CLIP model to accommodate new tasks by integrating a Mixture-of-Experts (MoE) adapter. Specifically, MoE consists of several LoRA adapter experts and routers, where the router calculates gating weights and uses the $TopK$ function to select the $k$ most relevant experts for learning the current task. To maintain the zero-shot recognition capabilities of the visual-language model, a Distribution Discriminative Automatic Selector (DDAS) is further introduced, which can automatically route in-distribution and out-of-distribution inputs to the MoE adapters and the original CLIP, respectively. Furthermore, the MoE-Adapters4CL framework has also been adapted for use in the CIL setting.

### v2

@@ 5.1.3. VLMs-based DIL
Vision-language models (VLMs) excel in domain-incremental learning, particularly through methods like S-Prompt (Wang et al., 2022b) (Figure 5(c)), which independently learns prompts across diverse domains using pre-trained VLMs. This approach introduces innovative image and language-image prompt acquisition techniques, utilizing a unified cross-entropy loss during training and a k-nearest neighbors (K-NN) method to identify domains during inference. In visual question answering, VQACL (Zhang et al., 2023d) enhances multimodal tasks through a dual-level task sequence. This framework uses a compositionality test to assess generalization to new skill-concept combinations and employs a representation learning strategy that differentiates sample-specific (SS) features, capturing unique input attributes, from sample-invariant (SI) features, which retain essential characteristics. DIKI (Tang et al., 2024) preserves pre-trained knowledge by using a residual attention mechanism to inject new information into a frozen backbone, minimizing interference. A distribution-aware calibration scheme ensures smooth integration of new knowledge while maintaining the VLM’s zero-shot capabilities. SC-MLLM (Liu et al., 2024) is a robotic manipulation framework that utilizes an exponential moving average (EMA) for continuous learning, reducing the risk of catastrophic forgetting.
5.2. Task-Incremental Learning
5.2.1. PLMs-based TIL
Traditional Methods.
Drawing inspiration from neurobiological mechanisms, HMI (Maekawa et al., 2023) (Figure 6(a)) integrates compressed representations of prior training instances, providing selective guidance for generating training samples. Unlike traditional Continual Relation Learning (CRL) that rely on large labeled datasets. Continual Few-Shot Relation Learning (CFRL) focuses on learning new relational patterns with minimal labeled data. ERDA (Qin and Joty, 2022a) tackles this challenge by selecting informative samples from unlabeled data and enforcing relational constraints in the embedding space. As for Lifelong Relation Extraction, EMR (Wang et al., 2019b) utilizes a working memory mechanism to selectively replay stored samples, helping the model retain prior knowledge while learning new tasks.
Learning user representations is fundamental for personalized recommender systems, but current approaches often train separate models for each task, which leads to inefficiency. To improve this, Conure (Yuan et al., 2021) introduces a framework that manages multiple tasks by pruning less critical parameters to accommodate new, task-specific ones. This method allows for positive transfer learning and preserves essential parameters to avoid knowledge loss. However, learning task-specific user representations for each task is impractical. Recent studies focus on universal user representations, which capture generalized user traits relevant to multiple tasks. TERACON (Kim et al., 2023) addresses this challenge by employing task-specific soft masks to isolate parameters and ensure effective knowledge retention.
@@ 5.2.1. PLMs-based TIL
Given that PLMs often suffers catastrophic forgetting when applied to CL scenarios, CTR (Ke et al., 2021a) tackles this by utilizing innovative techniques such as CL-plugins and task masking, which significantly improve knowledge transfer while reducing forgetting. This approach proves particularly effective on challenging benchmarks, such as the 20News dataset, where task dissimilarity often causes severe forgetting in other models. RecAdam (Chen et al., 2020) extends the concepts of EWC by applying a regularization term to the pre-trained weights, utilizing an annealing coefficient to progressively incorporate the importance of prior knowledge.
The diversity of text distributions and intents across domains poses challenges for scaling user intent detection (UIC) in industrial applications. MeLL (Wang et al., 2021) proposes a solution by leveraging global and local memory networks to maintain cross-task representations, which allows the model to continually adapt to new tasks without significant loss of performance. MeLL also employs an LRU policy for efficient memory management and controls parameter growth, making it scalable for real-world applications.
Recent advancements in conversational AI focus on overcoming the limitations of traditional chatbots, which rely on static knowledge bases and extensive manual data annotation. Lifelong INteractive learning in Conversation (LINC) introduces a dynamic learning framework that mirrors human cognitive processes, enabling chatbots to integrate and utilize knowledge in real-time (Liu and Mazumder, 2021; Mazumder and Liu, 2024; Liu, 2020). Structured with an Interaction Module, Task Learner, and Knowledge Store, LINC enhances chatbots by improving real-time information extraction, handling erroneous inputs, and refining conversational skills, thereby boosting linguistic and interactive capabilities.
Continual Pre-training Methods.
Continual pre-training represents a paradigm where PLMs are progressively enhanced by assimilating new knowledge from expanding datasets. ERNIE 2.0 (Sun et al., 2020b) illustrates this approach by incrementally constructing pre-training tasks, enabling the model to capture increasingly complex lexical, syntactic, and semantic nuances. Unlike traditional fixed-task training, it employs a continual multi-task learning framework to enhance its capabilities. Advancing this field further, RecyclableTuning (Qin et al., 2023) introduces two strategies: initialization-based and distillation-based. The former builds on fine-tuned weights from existing PLMs, while the latter reuses outdated weights to preserve knowledge and improve efficiency in future models.
Parameter-Efficient Tuning Methods.
Expounding upon the crucial need for more efficacious knowledge integration, ACM (Zhang et al., 2022) dynamically adjusts transformer architectures and leverages pseudo-experience replay to enhance knowledge transfer. Continual Learning of Few-Shot Learners (CLIF) addresses a gap between rapid generalization methods (such as few-shot learning) and traditional continual learning approaches, which are not inherently designed for such generalization. BiHNet-Reg (Jin et al., 2021) provides a solution by generating task-specific adapter weights through bi-level task representations and incorporating regularization.
Instruction Tuning-based Methods.
PCLL (Zhao et al., 2022), designed for task-oriented dialogue systems, uses a conditional variational autoencoder with prompts to generate pseudo samples that capture task-specific distributions. It incorporates distillation to reduce noise in these samples, outperforming baselines like Experience Replay and HAT in intent detection and slot filling. PP (Razdaibiedina et al., 2023) mitigates catastrophic forgetting and enables forward transfer without data replay or excessive task-specific parameters. This method appends new soft prompts to previously learned ones, keeping the base model unchanged. It surpasses LPT5, IDBR, and MBPA++ on both standard and custom continual learning benchmarks. Furthermore, ConTinTin (Yin et al., 2022) provides a framework for sequentially mastering tasks through textual instructions, supporting both forward and backward transfer of knowledge. DYNAINST (Mok et al., 2023), designed for lifelong in-context instruction learning, enhances a PLM’s generalization by combining parameter regularization with experience replay. Its Dynamic Instruction Replay, using Dynamic Instance Selection (DIS) and Dynamic Task Selection (DTS), optimizes memory and computation by selectively replaying relevant instances and tasks.
5.2.2. LLMs-based TIL
Recent attention has focused on the convergence of LLMs with CL methods, exemplified by significant contributions (Wang et al., 2024b; PENG et al., 2024). Benefiting from vast corpora and advanced hardware infrastructure, LLMs showcase remarkable capabilities in language comprehension and generation. However, challenges arise in scenarios involving sequential tasks, where LLMs often exhibit a decline in performance known as catastrophic forgetting.
@@ 5.2.2. LLMs-based TIL
DynaMind (Du et al., 2023) (Figure 6(b)) stands as a pioneering framework that enhances LLM output precision through the integration of memory mechanisms and modular operators. The framework consists of three key components: a memory module that stores and updates learned knowledge, a modular operator that processes incoming data, and a continual learning (CL) module that dynamically adjusts LLM parameters as new knowledge is introduced. In parallel, JARe (PENG et al., 2024) incorporates Dynamic Task-related Knowledge Retrieval (DTKR) to enable adaptive adjustments of language models for specific downstream tasks. By leveraging task distributions within vector space, JARe aims to optimize the continual learning process, streamlining task-specific knowledge adaptation.
@@
Large language models (LLMs) face significant challenges that limit their practical applications, such as high computational and memory demands, and vulnerability to catastrophic forgetting. These issues highlight the need for more efficient and robust training and deployment strategies. ConPET (Song et al., 2023), designed for the continual adaptation of LLMs across diverse tasks, leverages parameter-efficient tuning (PET) strategies to enhance both efficiency and performance. ConPET operates in two modes: Static ConPET, optimized for moderate-scale tasks, and Dynamic ConPET, tailored to large-scale tasks. Additionally, ELM (Jang et al., 2023) introduces a compact expert adapter trained for each task, with a retrieval mechanism that selects the most suitable expert LLM for new tasks. Furthermore, O-LoRA (Wang et al., 2023a) preserves task-specific knowledge by using low-rank approximations of previous task gradients, while the orthogonality constraint ensures task-specific subspaces remain distinct, minimizing interference from new tasks. This method demonstrates robustness across various task sequences.
@@
Continual-T0 (Scialom et al., 2022b) incorporates rehearsal techniques alongside instruction tuning of LLMs, showing strong performance across 70 datasets. However, its generalization to underrepresented languages remains limited. In response, InstructAlign (Cahyawijaya et al., 2023) proposes aligning new languages with those previously learned, leveraging abundant linguistic resources to mitigate catastrophic forgetting. The key innovation lies in enhancing language adaptation for instruction-tuned LLMs, particularly by incorporating underrepresented languages. Unlike traditional methods focused on single-task learning, RoboCoder (Li et al., 2024) enables robots to tackle a wide range of increasingly complex tasks by continuously refining and updating action codes based on environmental feedback. EUREKA (Ma et al., 2024) operates by leveraging LLMs to generate reward functions from the environment’s source code without requiring pre-defined templates or task-specific prompts. Through its evolutionary approach and in-context learning capabilities, it continually refines the rewards to achieve superior performance across diverse RL tasks.
@@ 5.2.3. VLMs-based TIL
CTP (Zhu et al., 2023) employs topology preservation and momentum contrast to maintain consistent relationships within sample mini-batches across tasks, thereby preserving the distribution of prior embeddings. ZSCL (Zheng et al., 2023) addresses the challenge of zero-shot transfer degradation in VLMs by employing a label-free dataset for distillation in the feature space, alongside weight regularization in the parameter space. Additionally, ZSCL has been adapted for the CIL setting, expanding its applicability to a wider range of continual learning scenarios. SND (Yu et al., 2024a) tackles catastrophic forgetting and zero-shot performance degradation in VLMs by leveraging a dual-teacher model. The pre-trained VLM retains zero-shot generalization, while the fine-tuned VLM preserves task-specific knowledge. A teacher selection mechanism quantifies discrepancies between their feature representations. Similarly, AwoForget (Zheng et al., [n. d.]) employs a graph-based multi-modal proximity distillation framework to mitigate forgetting in VLMs. It constructs a cross-modal graph to represent relationships between image-text pairs and adopts a dual-teacher distillation approach for maintaining both zero-shot and task-specific knowledge. This method also extends its applicability to CIL.
Parameter-Efficient Methods.
MoE-Adapters4CL (Yu et al., 2024b) introduces a parameter-efficient continual learning method to mitigate long-term forgetting in incremental learning of VLMs. Their approach involves dynamically extending the pre-trained CLIP model to accommodate new tasks by integrating a Mixture-of-Experts (MoE) adapter. Specifically, MoE consists of several LoRA adapter experts and routers, where the router calculates gating weights and uses the $TopK$ function to select the $k$ most relevant experts for learning the current task. To maintain the zero-shot recognition capabilities of the visual-language model, a Distribution Discriminative Automatic Selector (DDAS) is further introduced, which can automatically route in-distribution and out-of-distribution inputs to the MoE adapters and the original CLIP, respectively. Furthermore, the MoE-Adapters4CL framework has also been adapted for use in the CIL setting.
@@
By decoupling prompts and prompt interaction strategies, TRIPLET (Qian et al., 2023) (Figure 6(c)) effectively captures complex interactions between modalities. This includes specific designs for visual, textual, and fused prompts, as well as how to interact between different tasks through these prompts and retain crucial information, thereby reducing catastrophic forgetting. Decoupled prompts are designed to separate prompts in terms of multi-modality, layer-wise, and complementary, with each type of prompt containing learnable parameters intended to capture modality-specific knowledge from pre-trained VLMs and training data.